Cliff Corp. (CC) has an enterprise value (MV equity + debt – cash) of $330 million, $50 million in cash and $30 million in debt. CC has 10 million shares outstanding and is considering using its $50 million in cash to repurchase shares. What s CC’s share price prior to the prior to the repurchase? How many shares will CC repurchase? Suppose news is released that increases enterprise value to $350 million immediately after the repurchase, then what is CC’s new share price? If management knows about the news prior to the repurchase, should they repurchase before or after the news is released?
Answer #1

Solution:

Given:

Enterprise Value (EV) = $330 million

Cash = $50 million

Debt = $30 million

  • EV = MV of Equity + Debt - Cash

330 = Mv of Equity + 30 - 50

330 -30 + 50 = MV of Equity

It Implies:

MV of Equity = $350 million

Total Number of Shares Outstanding before repurchase = 10 million shares

  • MV of Equity Share before repurchase = 350/10

= $35 per share

As the current market value of shares are $35 per share it is assumed that it will be pucrhased at $35 per share from the open market.

  • Number of shares repurchased for cash = 50000000/35

= 1428571.429 i.e., 1428571 as part purchase of shares is not possible

  • Cash Available after repurchase = 50000000 - (14285571*35)

= 50000000 - 49999985

= $15

  • Calculation of new share price:

350,000,000 = MV of Equity + 30,000,000 -15

350,000,000 - 30,000,000 + 15 = MV of Equity

It implies:

MV of Equity after repurchase = 320000015

Number of Shares outstanding after repurchase = 10000000 - 14285571

= 8571429

MV of per share after repurchase = 320000015/8571429

= $37.33

After the news market value per share increases by $2.33 per share. Hence, it is right decision to repurchase the share at cheaper price. It results in more extinguishment of shares from open market at cheaper cost.